# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

### v1

@@ 4.2. Discussion on the general case
In general, the moduli space $M_{\beta,\chi}$ can be reducible even if $\beta$ is sufficiently positive; see [34, §5]. Moreover, even if we assume that $M_{\beta,\chi}$ is irreducible, it is usually hard to compute the codimension $c$ in Theorem 1.1. It is not necessarily true that $c$ is the same as $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$. In some nice cases where $c=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, the asymptotic behavior of $\operatorname{IH}^{k}(M_{\beta,\chi})$ can be described more precisely, provided that $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, or equivalently $N(\beta)$, can be calculated.
@@
Let $D_{1},\cdots,D_{\rho}$ be an orthogonal basis for $\operatorname{Num}(S)_{\mathbb{Q}}$, where $D_{1}$ is the class of an ample divisor. By Hodge index theorem, $D_{\ell}^{2}<0$ for $1<\ell\leq\rho$. Given any $C_{1}$, $C_{2}$ as in the statement, write
@@
for some $a_{j},a_{i,j}\in{\mathbb{Q}}$ ($i=1,2$), where $\equiv$ denotes the numerical equivalence. For $1<\ell\leq\rho$, let $n_{\ell}>0$ be the integer such that $A_{\ell}\vcentcolon=n_{\ell}D_{1}+D_{\ell}$ is ample. Then $D_{1}\cdot C_{i}>0$ and $A_{\ell}\cdot C_{i}>0$, which implies
@@
By the calculations in Section 5, we ask the following question.
Question 4.6.
@@
5. Stabilization in the Enriques and bielliptic cases
In this section, let $S$ be an Enriques or bielliptic surface. In particular, its canonical divisor $K_{S}$ is numerically trivial. When $\beta$ is a base-point-free, ample divisor on $S$, $|\beta|$ contains a connected smooth curve by Bertini’s theorem. Denote by $|\beta|^{\mathrm{sm}}\subset|\beta|$ the open subscheme parametrizing smooth curves.
To apply Theorem 1.1, we first prove the irreducibility of the moduli spaces of one-dimensional sheaves on these surfaces.
@@ 5. Stabilization in the Enriques and bielliptic cases
Suppose that $S$ is an Enriques or bielliptic surface, and $\beta$ is a base-point-free, ample divisor satisfying $\operatorname{gcd}(\beta\cdot H,\chi)=1$. Then $M_{\beta,\chi}$ is irreducible.
Proof.
Since the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ is irreducible and contains the smooth open subscheme $h^{-1}(|\beta|^{\mathrm{sm}})$ by [34, Proposition 2.2], there is a unique irreducible component $Y\subset M_{\beta,\chi}$ containing $h^{-1}(|\beta|^{\mathrm{int}})$. We prove the irreducibility of $M_{\beta,\chi}$ by contradiction. If $Y\not=M_{\beta,\chi}$, then take another irreducible component $Y^{\prime}\neq Y$. It follows from $Y^{\prime}\cap h^{-1}(|\beta|^{\mathrm{sm}})=\emptyset$ that $Y^{\prime}\subset h^{-1}(|\beta|\setminus|\beta|^{\mathrm{sm}})$. By [36, Corollary 1.3] and the fact that $K_{S}$ is numerically trivial, the fibers of $h:M_{\beta,\chi}\to|\beta|$ have the same dimension $p_{a}(\beta)$; therefore
(5.17) $\dim Y^{\prime}\leq\dim h^{-1}(|\beta|\setminus|\beta|^{\mathrm{sm}})<\dim|\beta|+p_{a}(\beta)=\beta^{2}+\chi({\mathcal{O}}_{S}),$
where the last equality follows from the Kodaira vanishing theorem and the Riemann–Roch formula. However, using the morphism $\operatorname{det}:M\to\operatorname{Pic}(X)$ in [11, Theorem 4.5.4]  [^2] Although [11, Theorem 4.5.4] is stated for moduli spaces of positive rank sheaves, the proof of $\dim_{[F]}M({\mathcal{Q}})\geq\dim_{[F]}M-\dim_{[\operatorname{det}F]}\operatorname{Pic}(X)$ does not rely on the positive rank assumption. and the Riemann–Roch formula, for any point ${\mathcal{F}}\in Y^{\prime}$, we have
$\dim Y^{\prime}$ $\geq\operatorname{ext}^{1}({\mathcal{F}},{\mathcal{F}})-\operatorname{ext}^{2}({\mathcal{F}},{\mathcal{F}})-h^{1}({\mathcal{O}}_{S})$
$=\beta^{2}+\chi({\mathcal{O}}_{S}),$
which contradicts (5.17). ∎
Then we assume the following property of $N(\beta)$, whose proof will be given in §5.1.2 and §5.2.2 by analyzing the Enriques and bielliptic cases separately.
Proposition 5.2.
If $\beta$ is an ample divisor on an Enriques or bielliptic surface, then
@@
Combining the above results with Theorem 1.1, we can prove the stabilization of intersection Betti numbers of $M_{\beta}=M_{\beta,1}$.
Theorem 5.3.
Suppose that $S$ is an Enriques or bielliptic surface. Fix $k\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on $S$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
(5.18) $\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all integer }d\geq d(\beta_{0},k),$
where $b_{k}^{\infty}$ is defined by (1.1).
Proof.
@@
    $d\beta_{0}$ is $\max\{2,2k-2\}$-very ample;
@@
    $2\dim|d\beta_{0}|\geq 3k$.
Since the fibers of $h:M_{d\beta_{0}}\to|d\beta_{0}|$ have the same dimension by [36, Corollary 1.3], we have
(5.19) $\operatorname{codim}(M_{d\beta_{0}}\setminus h^{-1}(|d\beta_{0}|^{\mathrm{int}}),M_{d\beta_{0}})$ $=\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$
@@
Hence, the result follows from Proposition 5.1 and Theorem 1.1 by taking $\beta=d\beta_{0}$. ∎
Remark 5.4.
When $\beta\cdot H$ and $\chi$ are not necessarily coprime, there can be strictly semistable sheaves in $M_{\beta,\chi}$ and thus $h:M_{\beta,\chi}\to|\beta|$ may not have constant fiber dimension by [36]. In this case, one could possibly prove the irreducibility of $M_{\beta,\chi}$ by showing that the corresponding moduli stack is irreducible (cf. [19, §2.6]), but the first equality of (5.19) may fail. Therefore, to ensure the equality in (5.18) holds for $M_{d\beta_{0},\chi}$, we need an additional assumption that $d\beta_{0}\cdot H$ is coprime with $\chi$.
@@ 5.1.1. Basic properties
Recall that an Enriques surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=h^{2}({\mathcal{O}}_{S})=0$. Its second Betti number is $b_{2}(S)=10$. The following property of linear systems on $S$ is well-known (e.g., [13, Theorem]).
@@
Let $D$ be a nonzero, nef, effective divisor on an Enriques surface. We have the following properties:
@@
    If $D^{2}=0$, then $D$ is numerically equivalent to $kE$ for some $k\in\mathbb{Z}_{>0}$ and a primitive, effective divisor $E$ with $E^{2}=0$. Moreover,
    $\dim|D|=\begin{cases}\lfloor\frac{k}{2}\rfloor&(\text{if }D=kE)\\ \lfloor\frac{k-1}{2}\rfloor&(\text{if }D=kE+K_{S}).\end{cases}$
@@
Any non-integral curves in $|d\beta|$ are in the image of $|C_{1}|\times|C_{2}|$ for finitely many decompositions $d\beta=C_{1}+C_{2}$ (up to linear or numerical equivalence) by Lemma 4.5, where $C_{1}$ and $C_{2}$ are nonzero effective divisors on $S$.
@@
Case 1.2: $C_{1}^{2}=0$ and $C_{2}^{2}>0$. By Lemma 5.5 (ii), $C_{1}=sE_{1}$ or $sE_{1}+K_{S}$ for some $s\in\mathbb{Z}_{>0}$ and a primitive effective divisor $E_{1}$ with $E_{1}^{2}=0$. Moreover,
(5.21) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{s}{2}\rfloor$
$\geq s\left(d\beta\cdot E_{1}-\frac{1}{2}\right)\geq d-\frac{1}{2}.$
Case 1.3: $C_{1}^{2}=C_{2}^{2}=0$. Let $s$ and $E_{1}$ be as in Case 1.2. By Lemma 5.5 (ii), $C_{2}=tE_{2}$ or $tE_{2}+K_{S}$ for some $t\in\mathbb{Z}_{>0}$ and a primitive effective divisor $E_{2}$ with $E_{2}^{2}=0$. Then we have $C_{1}\cdot C_{2}=stE_{1}\cdot E_{2}=d^{2}\beta^{2}/2>0$ and
(5.22) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{s}{2}\rfloor-\lfloor\frac{t}{2}\rfloor\geq\frac{d^{2}\beta^{2}-(s+t)}{2}$
@@
Case 2.2: $C_{1}^{2}=0$. Let $s$ and $E_{1}$ be as in Case 1.2. By Lemma 5.5 (ii),
(5.24) $\dim|d\beta|-\dim|C_{1}|\geq\frac{d^{2}\beta^{2}}{2}-\lfloor\frac{s}{2}\rfloor\geq\frac{s(d\beta\cdot E_{1}-1)+d\beta\cdot C_{2}}{2}\geq d-\frac{1}{2}.$
@@ 5.1.3. Refined invariants
Assume that $S$ is a generic Enriques surface and that the class of $\beta$ in $\operatorname{Num}(S)$ is not divisible by $2$ so that the moduli space $M_{\beta}=M_{\beta,1}$ is a smooth projective Calabi–Yau variety by [30]. Denote by $n_{\beta}^{i,j}$ the number defined by (2.4) with $\phi$ taken to be $h:M_{\beta}\to|\beta|$.
In order to apply Theorem 4.1 to compute $n_{\beta}^{i,j}$, we need the following lemma which is parallel to [20, Corollary 5.3]. For the reader’s convenience, we provide a proof below.
@@
Let $\beta$ be a base-point-free, ample divisor on a generic Enriques surface $S$ such that $2\nmid\beta$ in $\operatorname{Num}(S)$. When $i+j\leq N(\beta)$, we have
@@
where $h^{\circ}=h_{|\beta|^{\mathrm{int}}}:h^{-1}(|\beta|^{\mathrm{int}})\to|\beta|^{\mathrm{int}}$.
@@
Let $W=M_{\beta}\setminus h^{-1}(|\beta|^{\mathrm{int}})$. Denote by $\iota:W\hookrightarrow M_{\beta}$ the closed embedding, and by $\jmath:h^{-1}(|\beta|^{\mathrm{int}})\hookrightarrow M_{\beta}$, $\jmath^{\circ}:|\beta|^{\mathrm{int}}\to|\beta|$ the open embeddings. Since the fibers of $h$ have the same dimension, we have
@@
Consider the exact triangle
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath_{*}\jmath^{*}{\mathbb{Q}}_{M_{\beta}}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}$
which yields another exact triangle
$h_{*}\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow h_{*}{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath^{\circ}_{*}h^{\circ}_{*}{\mathbb{Q}}_{h^{-1}(|\beta|^{\mathrm{int}})}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}.$
By the property of Verdier duals ([23, Proposition 5.3.9]) and the smoothness of $M_{\beta}$,
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\cong(D_{M_{\beta}}\iota_{*}{\mathbb{Q}}_{W})[-2\dim M_{\beta}],$
@@
$H^{k}(\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})=H^{2\dim M_{\beta}-k}(W)^{\vee}=0.$
By taking the long exact sequence of cohomologies, the vanishing of $H^{k}(\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})$ shows that for $\ell\leq N(\beta)$, the restriction map induces an isomorphism
@@
Since the restriction to $h^{-1}(|\beta|^{\mathrm{int}})$ clearly preserves a relatively ample class (relative to $h$), it follows from [6, Lemma 3.3] that $n_{\beta}^{i,j}\leq n_{h^{\circ}}^{i,j}$ for $i+j\leq N(\beta)$. Then the result follows by (4.12) and (5.25). ∎
Now we can prove the stabilization of $n_{\beta}^{i,j}$ for the moduli space of one-dimensional sheaves on a generic Enriques surface.
Theorem 5.7.
Fix $i,j\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$, $i$ and $j$) such that
(5.26) $n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j}\quad\text{for all integer $d\geq d(\beta_{0},i,j)$ with }2\nmid d,$
where $n_{\infty}^{i,j}$ is determined by
$\sum_{i,j\geq 0}n_{\infty}^{i,j}q^{i}t^{j}=(1-qt)\prod_{m\geq 1}\frac{1}{(1-q^{m+1}t^{m-1})(1-q^{m}t^{m})^{10}(1-q^{m-1}t^{m+1})}.$
Proof.
Choose the integer $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ such that for all $d\geq d(\beta_{0},i,j)$, the following properties hold:
@@
    $d\beta_{0}$ is $\max\{0,i-1\}$-very ample;
@@
Condition (3) is equivalent to $2\dim|d\beta_{0}|\geq 3i+j$ by Lemma 5.5 (i). By Lemma 5.6 and Theorem 4.1, taking $U=|d\beta_{0}|^{\mathrm{int}}$, we have
$n_{d\beta_{0}}^{i,j}=n_{h_{U}}^{i,j}=n_{\infty}^{i,j},$
which proves (5.26). ∎
@@
Fix $k\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
@@
where $b_{k}^{\infty}$ is determined by (1.1).
Remark 5.9.
By (5.20)–(5.22) and [35, Proposition 2.5], we can take
(5.27) $d(\beta_{0},i,j)=\max\left\{2,\,i+1,\,\left\lceil\frac{i+j+2}{2}\right\rceil,\,\left\lceil\frac{i+j+6}{2\sqrt{2\beta_{0}^{2}}}\right\rceil,\,\left\lceil\sqrt{\frac{2i+2j+6}{\beta_{0}^{2}}}\,\right\rceil\right\}$
in Theorem 1.4 (Theorem 5.7).

### v2

@@ 4.3. Discussion on the general case
In general, the moduli space $M_{\beta,\chi}$ can be reducible even if $\beta$ is sufficiently positive (see [42, §5]). Moreover, even if we assume that $M_{\beta,\chi}$ is irreducible, it might not be true that the codimension $c$ in Theorem 1.1 is the same as $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$. In the cases where $c=\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, the asymptotic behavior of $I\!H^{k}(M_{\beta,\chi})$ can be described more precisely, provided that $\operatorname{codim}(|\beta|\setminus|\beta|^{\mathrm{int}},|\beta|)$, or equivalently $N(\beta)$ (defined in §4.2), can be calculated.
@@
Let $D_{1},\cdots,D_{\rho}$ be an orthogonal basis for $\operatorname{Num}(S)_{\mathbb{Q}}\vcentcolon=\operatorname{Num}(S)\otimes_{\mathbb{Z}}{\mathbb{Q}}$, where $D_{1}$ is the class of an ample divisor. By the Hodge index theorem, $D_{\ell}^{2}<0$ for $1<\ell\leq\rho$. Given any $C_{1}$, $C_{2}$ as in the statement, write
@@
for some $a_{j},a_{i,j}\in{\mathbb{Q}}$, where $\equiv$ denotes the numerical equivalence. For $1<\ell\leq\rho$, let $n_{\ell}>0$ be an integer such that $A_{\ell}\vcentcolon=n_{\ell}D_{1}+D_{\ell}$ is ample. Then $D_{1}\cdot C_{i}>0$ and $A_{\ell}\cdot C_{i}>0$, which implies
@@
Motivated by the del Pezzo case ([36, Proposition 1.5]), along with the cases studied in Section 5, we ask the following question.
Question 4.10.
@@
5. Stabilization for minimal surfaces of Kodaira dimension zero
Throughout this section, except in §5.4, we consider a minimal surface $S$ of Kodaira dimension 0. In particular, its canonical divisor $K_{S}$ is numerically trivial. In this situation, the following proposition gives an affirmative answer to Question 4.10, which we then use to prove Theorem 1.4.
@@ 5. Stabilization for minimal surfaces of Kodaira dimension zero
If $\beta$ is an ample divisor on a minimal surface $S$ of Kodaira dimension 0, then
@@
The proof of Proposition 5.1 will be given in §5.1.2, §5.2.2, and §5.3, where we analyze the Enriques, bielliptic, K3 and Abelian cases separately.
Together with Proposition 2.2 and Proposition 5.1, Theorem 1.1 implies the stabilization of the intersection Betti numbers of $M_{\beta,\chi}$.
Proof of Theorem 1.4.
@@
    $d\beta_{0}$ is $\max\{1,2k-2\}$-very ample;
@@
    $2\dim|d\beta_{0}|=2\chi({\mathcal{O}}_{S}(d\beta_{0}))-2\geq 3k$.
By Proposition 2.2 (i), $M_{d\beta_{0},\chi}$ is irreducible, and by Proposition 2.2 (ii), we have
$\operatorname{codim}(M_{d\beta_{0},\chi}\setminus h^{-1}(|d\beta_{0}|^{\mathrm{int}}),M_{d\beta_{0},\chi})$ $=\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$
@@
Hence, the result follows from Theorem 1.1 with $\beta=d\beta_{0}$. ∎
@@ 5.1.1. Basic properties
Recall that an Enriques surface $S$ is a minimal surface of Kodaira dimension 0 with $h^{1}({\mathcal{O}}_{S})=h^{2}({\mathcal{O}}_{S})=0$. Its second Betti number is $b_{2}(S)=10$. The following property of linear systems on $S$ is well-known (e.g., [19, Theorem]) and was used in [37, Proposition 1.14] to compute the codimension of the locus of reducible curves in a linear system.
@@
Let $D$ be a nonzero effective nef divisor on an Enriques surface $S$.
@@
    If $D^{2}=0$, then $D$ is linearly equivalent to either $mE$ or $mE+K_{S}$ for some $m\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E$ with $E^{2}=0$. Moreover,
    $\dim|D|=\begin{cases}\lfloor\frac{m}{2}\rfloor&(\text{if }D\sim mE)\\ \lfloor\frac{m-1}{2}\rfloor&(\text{if }D\sim mE+K_{S}).\end{cases}$
@@
By Lemma 4.9, there are finitely many pairs of nonzero effective divisors $(C_{1},C_{2})$ (up to linear or numerical equivalence) with $d\beta=C_{1}+C_{2}$. Any non-integral curve in $|d\beta|$ lies in the image of $|C_{1}|\times|C_{2}|$ for one of these pairs.
@@
Case 1.2: $C_{1}^{2}=0$ and $C_{2}^{2}>0$. By Lemma 5.2 (ii), $C_{1}$ is linearly equivalent to $m_{1}E_{1}$ or $m_{1}E_{1}+K_{S}$ for some $m_{1}\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E_{1}$ with $E_{1}^{2}=0$. Moreover,
(5.19) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{m_{1}}{2}\rfloor$
$\geq m_{1}\left(d\beta\cdot E_{1}-\frac{1}{2}\right)\geq d-\frac{1}{2}.$
Case 1.3: $C_{1}^{2}=C_{2}^{2}=0$. Let $m_{1}$ and $E_{1}$ be as in Case 1.2. By Lemma 5.2 (ii), $C_{2}$ is linearly equivalent to $m_{2}E_{2}$ or $m_{2}E_{2}+K_{S}$ for some $m_{2}\in\mathbb{Z}_{>0}$ and some primitive effective divisor $E_{2}$ with $E_{2}^{2}=0$. Then we have $C_{1}\cdot C_{2}=m_{1}m_{2}E_{1}\cdot E_{2}=d^{2}\beta^{2}/2>0$ and
(5.20) $\dim|d\beta|-\dim(|C_{1}|\times|C_{2}|)$ $\geq C_{1}\cdot C_{2}-\lfloor\frac{m_{1}}{2}\rfloor-\lfloor\frac{m_{2}}{2}\rfloor\geq\frac{d^{2}\beta^{2}-(m_{1}+m_{2})}{2}$
@@
Case 2.2: $C_{1}^{2}=0$. Let $m_{1}$ and $E_{1}$ be as in Case 1.2. By Lemma 5.2 (ii),
(5.22) $\dim|d\beta|-\dim|C_{1}|\geq\frac{d^{2}\beta^{2}}{2}-\lfloor\frac{m_{1}}{2}\rfloor\geq\frac{m_{1}(d\beta\cdot E_{1}-1)+d\beta\cdot C_{2}}{2}\geq d-\frac{1}{2}.$
@@ 5.1.3. Refined invariants
Assume that $S$ is a generic Enriques surface. For an ample divisor $\beta$ whose class in $\operatorname{Num}(S)$ is not divisible by $2$, the moduli space $M_{\beta}=M_{\beta,1}$ is a smooth projective Calabi–Yau variety by [37]. The linear system $|\beta|$ contains a connected smooth curve (e.g., [37, Corollary 1.13]). Denote by $n_{\beta}^{i,j}$ the number defined by (2.6) with $\varphi$ taken to be $h:M_{\beta}\to|\beta|$.
In order to apply Theorem 4.5 to compute $n_{\beta}^{i,j}$, we need the following lemma which is parallel to [26, Corollary 5.3]. For the reader’s convenience, we provide a proof below.
@@
Let $\beta$ be an ample divisor on a generic Enriques surface $S$ with $2\nmid\beta$ in $\operatorname{Num}(S)$. For any $i,j\in\mathbb{Z}_{\geq 0}$ satisfying $i+j\leq N(\beta)$, we have
@@
where $h^{\circ}=h_{|\beta|^{\mathrm{int}}}:h^{-1}(|\beta|^{\mathrm{int}})\to|\beta|^{\mathrm{int}}$ is the restriction of $h:M_{\beta}\to|\beta|$.
@@
Let $W=M_{\beta}\setminus h^{-1}(|\beta|^{\mathrm{int}})$. Denote by $\iota:W\hookrightarrow M_{\beta}$ the closed embedding, and by $\jmath:h^{-1}(|\beta|^{\mathrm{int}})\hookrightarrow M_{\beta}$, $\jmath^{\circ}:|\beta|^{\mathrm{int}}\hookrightarrow|\beta|$ the open embeddings. By Proposition 2.2 (ii),
@@
Consider the exact triangle (all functors are derived)
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\longrightarrow{\mathbb{Q}}_{M_{\beta}}\longrightarrow\jmath_{*}\jmath^{*}{\mathbb{Q}}_{M_{\beta}}\stackrel{{\scriptstyle+1}}{{\longrightarrow}}.$
By the property of Verdier duals [30, Proposition 5.3.9] and the smoothness of $M_{\beta}$,
$\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}}\simeq(D_{M_{\beta}}\iota_{*}{\mathbb{Q}}_{W})[-2\dim M_{\beta}],$
@@
${\mathbb{H}}^{k}(M_{\beta},\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})\cong H^{2\dim M_{\beta}-k}(W)^{\vee}=0.$
By the long exact sequence in hypercohomology, the vanishing of ${\mathbb{H}}^{k}(M_{\beta},\iota_{*}\iota^{!}{\mathbb{Q}}_{M_{\beta}})$ shows that for $\ell\leq N(\beta)$, the restriction map induces an isomorphism
@@
Since the restriction to $h^{-1}(|\beta|^{\mathrm{int}})$ clearly preserves a relatively ample class (relative to $h$), it follows from [7, Lemma 3.3] that $n_{\beta}^{i,j}\leq n_{h^{\circ}}^{i,j}$ for $i+j\leq N(\beta)$. Then the result follows by (4.13) and (5.23). ∎
Now we can prove the stabilization of $n_{\beta}^{i,j}$ for moduli spaces of one-dimensional sheaves on a generic Enriques surface.
Proof of Theorem 1.6.
Choose $d(\beta_{0},i,j)$ such that for all $d\geq d(\beta_{0},i,j)$, the following properties hold:
@@
    $d\beta_{0}$ is (i1)(i-1)-very ample;
@@
Property (3) is equivalent to $2\dim|d\beta_{0}|\geq 3i+j$ by Lemma 5.2 (i). Applying Lemma 5.3 and Theorem 4.5 with $\beta=d\beta_{0}$ and $U=|d\beta_{0}|^{\mathrm{int}}$, we have
$n_{d\beta_{0}}^{i,j}=n_{h_{U}}^{i,j}=[H(q,t)]^{i,j}=n_{\infty}^{i,j},$
which completes the proof. ∎
@@
Given any ample divisor $\beta_{0}$ on a generic Enriques surface $S$ with $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$ and any $k\in\mathbb{Z}_{\geq 0}$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ such that
@@
where $b_{k}^{\infty}$ is defined by (1.1).
Remark $5.$
Note that a generic Enriques surface does not contain any (2)(-2)-curves. By (5.18)–(5.20) and [44, Proposition 2.5], we can take
(5.24) $d(\beta_{0},i,j)=\max\left\{i+1,\,\left\lceil\frac{i+j+2}{2}\right\rceil,\,\left\lceil\frac{i+j+6}{2\sqrt{2\beta_{0}^{2}}}\right\rceil,\,\left\lceil\sqrt{\frac{2i+2j+6}{\beta_{0}^{2}}}\,\right\rceil\right\}$
in Theorem 1.6.